How can I find the values of the constants of the system of equations?
6.5=936a+b
8.9=1256a+b

Answer by checkley79(3341)

6.5=936a+b
-(8.9=1256a+b) SUBTRACT
-----------------------
-2.4=-320a
a=-2.4/-320
a=.0075 ans.
6.5=936(.0075)+b
6.5=7.02+b
b=6.5-7.02
b=-.52 ans.
Proof:
8.9=1256(.0075)+-.52
8.9=9.42-.52
8.9=8.9